How many degrees is the measure of the supplement of an angle whose complement measures 50 degrees?
Answer by gonzo(654)

x = angle
90-x = complement of angle.
180-x = supplement of angle.
----
90-x = 50 (given)
x = 40
180-40 = 140
-----
angle = 40
complement of angle = 50 (40 + 50 = 90)
supplement of angle = 140 (40 + 140 = 180)
-----